Adv. Math. Econ. 8, 185-214 (2006)

On extensive form implementation of equilibria in differential information economies

Dionysius Glycopantis^{1}, Allan Muir^{2} and Nicholas C. Yannelis^{3}

1 Department of Economics,City University,Northamton Square,London EC1V0HB,UK

2 Department of Economics,City University,Northamton Square,London EC1V0HB,UK

3 Department of Economics,University of Illinois at Urbana-Champaign,IL 61820,USA

Received: May 16, 2005 Revised: September 12, 2005

JEL classification: D5, D82, C71, C72.

Abstract. We investigate the possibility of a dynamic explanation of the equilibrium ideas in terms of the perfect Bayesian equilibrium (PBE) (or sequential equilibrium). In particular, we take an equilibrium outcome which has been found by means of a static optimizing behavior and ask the following question. Can this outcome be supported, (or implemented), as a PBE of an extensive form game of a reasonable form? We provide a positive answer for solution concepts which are incentive compatible and a negative one for those which are not.

Key words: differential information economy, Walrasian expectations equilibria, rational expectations equilibria, weak fine core, private core, weak fine value, private value, coalitional Bayesian incentive compatibility, game trees, perfect Bayesian equilibrium